5. (8 points) A force 250 with the positive x axis. Draw the force vector on the x-y plane scalar componen vector has a magnitude of 15.0 newtons and makes at an angle of t and (b) the y scalar component of the vector and find (a) the x 6. (8 pts.) A force vector has x component of -41 newtons and a y component of -27 newtons. Find (a) the magnitude and (b) the direction the force vector.
Answer #1

5) angle that vector makes with positive x axis=250 degrees counterclockwise

Therefore, the vector makes an angle of 250-180=70 drgrees with negative x axis counterclockwise

a)X component of vector=-15*cos70=-5.13

b)Y component of vector=-15*sin70=-14.10

6) magnitude of force=sqrt(41²+27²)=49.09

Angle vector makes with positive x axis=180+arctan(27/41)=213.3 degress counterclockwise